Your friend is asking you to invest in his bakery, and you have 35000. The returns are volatile, and you may get either 43000 with a probability of 0.44 or 27000 with a probability of 0.56.


1
Expert's answer
2021-10-05T11:02:30-0400

Those are the questions:

a. What is the expected value of the investment?  


b. Will you buy the refrigerator? Why? 


c. Identify the mean, variance, and standard deviation.


(a) The expected value is "43000*0.44+35000*0.56=34040"

(b) It would more likely unprofitable to buy refrigerator cause there is higher probability of disadvantageously outcome

(c) The mean is "(43000+27000)*0.5=35000"

The variance is "43000^{2}*0.44+27000^{2}*0.56 - 34040^{2}=63078400"

The standart deviation is the square root from the variance and equal to 7942.19